write down the joint probability density function of the bivariate normal distribution for the random variables x1 and x2, defining all parameters


1
Expert's answer
2021-10-04T19:38:12-0400

"f_{X_1X_2}(x_1.x_2)=\frac{1}{2\pi \sigma_{X_1}\sigma_{X_2}\sqrt{1-\rho^2}}\cdot exp\{-\frac{1}{2(1-\rho^2)}[(\frac{x_1-\mu_{X_1}}{\sigma_{X_1}})^2+(\frac{x_2-\mu_{X_2}}{\sigma_{X_2}})^2-"


"-2\rho \frac{(x-\mu_{X_1})(x_2-\mu_{X_2})}{\sigma_{X_1}\sigma_{X_2}}]\}"


where "\mu_{X_1},\ \mu_{X_2}" are means,

"\sigma_{X_1},\ \sigma_{X_1}" are standard deviations,

"\rho" is correlation coefficient.